Računalništvo I

Spisek študentov

17. februar 1998

• Ahčin Tatjana

• Kerčmar Nina

• Kosec Jana

• Kosta Maja

• Lunar Klavdija

• Majerhold Tomaž

• Mastnak Matjaž

• Merc Urška

• Nežmah Barbara

• Perčič Simona

• Rici Jurij

• Rojec Aljoša

• Šabič Daliborko

• Verdinek Irena

• Žabkar Jure Če obiskuje predmet še kakšen študent, naj me o tem nemudoma obvesti. Andrej